• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 15:59:55 on November 15, 1998

    Entry number 5487

    keyword=Beam Time Accounting
    Hall A

    11/15/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: S. Incerti
    Shift Workers: F. Sabatie, J. Calarco, K. Kramer

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    1 1 0.25 1 1 0.5 1 0.67 6.42
    Beam Available
    but Not in Use
    0 0 0 0 0 0.5 0 0.33 0.83
    Beam Not Available
    or Unacceptable
    0 0 0.75 0 0 0 0 0 0.75
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 1 1 1 0.75 7.75
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0.25 0.25
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: - smooth running perpendicular configuration, Pe= 3.631 and Ph=2.433 - Bscope at 9:45 AM - loss of end of run 2563 at 10:25 ? - back at 11:00 - raster checks and harp scans starting at 13:10 - NMR polarization gives 29 to 30% - back to transverse configuration at 14:30 - switched to parallel configuration at15;30 - new momentum on electron arm : Pe=3.0858 GeV - resume data taking at 15:50

    Data Entry by: incerti